# A Survey of Conversational Search

### v1

@@ Abstract
Conversational Search; Query Reformulation; Search Clarification; Conversational Retrieval and Generation
@@ 1 Introduction
Search engines have become integral to modern society, serving as essential tools for fulfilling users’ information needs. Their development has been significantly accelerated by advances in artificial intelligence (AI) [1]. Recently, with the rapid advancements in natural language processing (NLP) technologies, particularly the emergence of large language models (LLMs) [2], search engines have evolved to offer more intelligent and interactive user experiences. A notable advancement in this domain is conversational search, an emerging paradigm that provides natural language interactions for complex and accurate information access. Commercial conversational AI search engines, such as Perplexity.ai and SearchGPT, [^1] Perplexity.ai: https://www.perplexity.ai/, SearchGPT: https://searchgpt.com/, have already been deployed, rapidly attracting a large user base and keeping significant growth.
@@
Compared to traditional search engines that rely on keywords or short phrases, conversational search leverages natural language dialogue for interactions [3, 4, 5, 6, 7], greatly enhancing the efficiency of information exchange and optimizing the user experience. This approach supports more complex user queries, manages longer and more intricate historical contexts, and offers comprehensive capabilities for information integration and processing. Moreover, conversational search enables diverse methods of information delivery and possesses the ability to engage in active interactions with users [8, 9, 10]. For example, these systems can proactively ask clarifying questions [11, 12, 13], make recommendations [14], guide users in better expressing their needs through dialogue. Figure 1 illustrates the differences between traditional search and conversational search. In this real example, a traditional search engine like Google fails to maintain context between queries. After the user queries, “What is information retrieval”, and then follows up with “Tell me some of its famous scholars”, the search engine provides irrelevant results because it does not recognize that the user is still referring to “information retrieval scholars”. In contrast, the conversational search engine Perplexity.ai successfully manages multi-turn queries by retaining context, correctly interpreting the follow-up question, and providing accurate information about notable scholars in the filed of information retrieval.
@@
Figure 1: Comparison between a traditional search engine Google and a new conversational search engine Perplexity.ai. The term “its” in the second query refers to the term “information retrieval” in the first query. However, it is challenging for traditional search engines to understand such omissions in user queries.
A typical conversational search system comprises several essential modules, each involving a variety of advanced technologies. As this field continues to evolve, it presents both significant opportunities and challenges. In this survey, we will explore and analyze the most critical components that constitute a conversational search system, following the entire information flow. Our focus will include query reformulation, search clarification, conversational retrieval, and response generation. These components are integrated to work collaboratively, enabling more natural and intuitive interactions between users and the system, ultimately enhancing the overall search experience.
@@
Search Clarification. Search clarification is another crucial component in conversational search systems. It enables users to refine their search queries through interactive dialogue, which can cover scenarios like seeking information, performing tasks, or navigating websites [19, 20, 21]. Users often express their search intents ambiguously or in multifaceted ways. To improve understanding, the system can pose clarifying questions, such as “Do you mean [specific term]?” or “Can you provide more details about [topic]?”, rather than directly answering the query. When the system detects that clarification is needed, it proactively asks these questions to better understand the user’s intent. This approach ensures that the system provides more accurate and relevant search results, enhancing the overall user experience by making the search process more personalized and effective. The primary challenges in search clarification include accurately and efficiently identifying when clarification is necessary and generating appropriate clarifying questions.
@@
Our paper differentiates itself from prior surveys on conversational search [54, 55, 56, 57] in two significant ways. First, earlier surveys were conducted prior to the emergence of ChatGPT [^2] ChatGPT: https://openai.com/chatgpt/ and primarily focused on traditional methods without LLMs. LLMs have revolutionized the techniques and application scenarios in this area, leading to the development of real commercial conversational search engines. Our survey incorporates the latest LLM techniques in conversational search and discusses the associated new challenges and opportunities. Second, our survey adopts a more systematic approach. We structure our review from the perspective of information flow in conversational search engines, focusing specifically on the four key modules (i.e., query reformulation, clarification, conversational retrieval, and conversational response generation). A broad overview of conversational search is presented in Figure 3, which categorizes existing studies using a structured taxonomy.
@@
Assessment (Sec 2.3) [ Dataset Analysis: CANARD [75], QReCC [76], TREC-CAsT [77, 78, 79, 80], leaf, text width=27.8em ] [ Assessment: Lexical Overlapping [75, 76], End-to-End [62, 6], leaf, text width=27.8em ] ] ] [ Search Clarification
@@
Tree-of-clarification [113], CAmbigNQ [20], leaf, text width=27.8em ] ] [ Search Clarification
@@
(Sec 4.2)[ Implicit: ConvDR [122], COTED [24], ZeCo [123], QRACDR [124], leaf, text width=27.8em ] [ Explicit: ConvRelExpand [62], HAConvDR [27], leaf, text width=27.8em ] ] [ Data Augmentation
@@
Results (Sec 5.1)[ Fine-Grained RAG [136], Conv-CoA [121], leaf, text width=27.8em ] ] [ Context Dependency
@@
Attribution (Sec 5.3)[ WebGPT [140], LaMDA [141], WebBrain [142] , leaf, text width=27.8em ] ] ] [ Domain-specific and
@@
LAPS [154], PersonalizedCIR [155], ConvWebAgent [156], leaf, text width=27.8em ] ] ] [ Benchmark and
@@
SIPCIS [168], PACS [169],leaf, text width=27.8em ] [ Generation-based: eRAG [170], PRGR [171], EvalLLMQA [172] ,leaf, text width=27.8em ] ] [ Benchmark (Sec 7)[ MSDialog [173], TREC CAsT-19 [77], Qulac [19], CANARD [75],
@@
FOCUSQA [186], TREC iKAT-23 [10], ConvRAG [136],
@@ 2.1 Query Reformulation in Conversational Search
In conversational search, query reformulation is critical due to the complex nature of user intent, which is often obscured within follow-up questions and results in context-dependent queries [77, 78, 80]. Unlike traditional information retrieval, conversational search must address two key semantic phenomena: Anaphora and Ellipsis [65, 70]. Anaphora involves expressions of which semantics rely on previous context, while Ellipsis refers to the omission of semantic elements understood from prior context. These phenomena necessitate that current queries are semantically linked to previous historical context, e.g., query-answer pairs, making context tracking essential for accurate responses. Figure 4 shows a case of query reformulation in conversational search. Adaptive techniques [75] are required to resolve references and fill in omitted information based on conversation history. Consequently, effective query reformulation must be dynamic and context-aware, continuously adapting to the evolving discourse. From the literature, the existing methods can be divided into three categories: 1) Query expansion, 2) Query rewriting, and 3) Hybrid, which are introduced as follows.
Figure 4: An example illustrating the semantic phenomena of Anaphora and Ellipsis in conversational search. Query reformulation plays a crucial role in addressing these semantic gaps.
@@ 2.1.1 Query Expansion
Relevance term classification transforms query reformulation as a binary classification task where terms in the query context are labeled as relevant or irrelevant. The relevant terms are then appended to the original query as expansion. This approach systematically incorporates important terms, thereby improving the retrieval performance. Typically, Kumar et al. [58] develop a binary term classifier with weak-supervision signals to select relevant terms from the dialog history and pseudo relevance feedback (PRF) to de-contextualize the conversational query. Voskarides et al. [59] propose QuReTeC with distant supervision, whose pseudo labels are readily available and can be inferred from user-system interactions. Lin et al. [60, 195] estimate term importance during query reformulation, which evaluates the significance of different terms within the context, ensuring that the most critical terms are retained or emphasized in the reformulated query.
Besides, heuristic methods focus on leveraging linguistic features to reformulate queries, which leverage techniques like part-of-speech tagging, dependency parsing, and co-reference resolution to reconstruct the original query by identifying grammatical structures and resolving references to the previous context. By utilizing syntactic and semantic analysis, these methods [61, 59] enhance the clarity and relevance with the context of the reformulated query and thus improve the search results.
@@ 2.2 Analysis of Existing Datasets
Typical query reformulation datasets designed for conversational search include CANARD [75], QReCC [76], the TREC Conversational Assistance Track (CAsT) series [77, 78, 79, 80] etc. Their construction involves both manual annotation and automatic generation. As aforementioned discussion, the quality of the annotation would greatly affect the search performance. Though with high cost, the human annotators manually rewrite the original queries to provide clear and explicit context, ensuring high-quality annotation. This is done in the TREC Conversational Assistance Track (CAsT) series across the test set and exemplified by the CANARD dataset in training data. The rewritten queries provided in the QReCC dataset are inherited from CANARD and expanded automatically by replacing the pronouns with anaphora resolution rules. As a result, its linguistic phenomenon is not as abundant as the fully manually constructed datasets. Besides, the relevance judgments for the documents associated with the original question are determined by whether containing the answers from the annotators, which might not guarantee the corresponding rewritten queries are optimal search queries. Thus, the discrepancy between document assessment and rewritten query annotation for each query turn should be carefully addressed when designing a query reformulation model for downstream search tasks.
@@ 2.4 Limitation and Discussion
Paradigm of Query Reformulation in Future. The goal of query reformulation in existing retrieval systems, either ad-hoc search or conversational search, is to supply missing search aspects or address ambiguity in the original query. With the development of large language models, some new paradigms of query reformulation have emerged, e.g., collaboration with different search agents [198], proactively providing reformulated query [199], reformulating query with reasoning chain and generative language models [200, 201, 202]. These studies improve query reformulation with various aspects and target different goals without a well-defined paradigm. It is unclear what is the destination of query reformulation under the evolved conversational search paradigm, which requires further exploration from the community.
@@ 3 Search Clarification
Conversational search provides users with an interactive interface enabling them to seek information, conduct behavior, or find a website with multi-turn interactions. During the conversation, users may deliver ambiguous, faceted, or unspecific search intents, which should be clarified by the systems for exact query understanding. To this end, the systems are expected to actively ask the users clarifying questions to help them re-articulate their intents. In this section, we summarize studies about asking clarifying questions in conversational search systems. Since the scenarios of different conversational search systems might be various, we first categorize the clarification scenarios as four main parts including (1) Clarification for conversational document retrieval, (2) Web search clarification, (3) Search clarification for question answering, and (4) Domain-specific search clarification. As for the domain-specific search clarification, we further divide it into different specific domains including Conversational Recommender System, legal search clarification, search clarification for programming, etc. We further summarize studies focusing on search clarification with Large Language Models (LLMs) and finally discuss future directions. In each scenario, the system may behave differently as shown in Figure 5.
@@ 3.1 Clarification for Conversational Retrieval
Conversational retrieval is a crucial part of a conversational search system. In this retrieval process, search clarification is defined as the multi-turn interaction between the user and the system [19] as shown in Figure 5 (a). Formally, in each interaction turn $k$, the user inputs a query ($k=1$) or responds to the clarifying question returned by the system ($k>1$). For the system, it asks a clarifying question and retrieves relevant documents using obtained contexts simultaneously in each turn. The fundamental idea of the conversational retrieval system with clarification is to guide users to deliver more useful information through multiple turns of querying so that better documents can be returned by the system. Thus, adding clarification information to the retrieval model can enable the system to retrieve more relevant documents that are in line with the user’s intentions [203]. In this section, we first summarize the available datasets for clarification in conversational retrieval and then introduce clarifying question selection and generation methods in this setting. We further explore analytical studies that focus on data features instead of proposing new methods. Finally, we introduce some other relevant studies.

### v2

@@ Abstract
Conversational Search, Query Reformulation, Search Clarification, Conversational Retrieval and Generation, Domain-specific and User-centric Application, Benchmark and Evaluation
@@ 1 Introduction
Search engines have become integral to modern society, serving as essential tools for fulfilling users’ information needs. Their development has been significantly accelerated by advances in artificial intelligence (AI) [1, 2]. Recently, with the rapid advancements in natural language processing (NLP) technologies, particularly the emergence of large language models (LLMs) [3], search engines have evolved to offer more intelligent and interactive user experiences. A notable advancement in this domain is conversational search, an emerging paradigm that provides natural language interactions for complex and accurate information access. Commercial conversational AI search engines, such as Perplexity.ai and SearchGPT, [^1] Perplexity.ai: https://www.perplexity.ai/, SearchGPT: https://searchgpt.com/, have already been deployed, rapidly attracting a large user base and keeping significant growth. Conceptually, conversational search represents a paradigm shift towards interactive information access, which moves beyond single-shot queries to enable users to fulfill complex, often evolving, information needs through multi-turn natural language dialogue with a system. The core functionality of conversational search is to support information goal evolution, where a user’s initial, possibly vague, need is refined or shifted through interaction. This necessitates a shared contextual understanding built turn-by-turn and often involves mixed initiative, where both the user and the system can guide the dialogue, e.g., the user asks or the system seeks clarification and proactively suggests refinements. Ultimately, the search goal extends beyond mere relevance to encompass a richer user experience, focusing on efficiency, reducing cognitive load, enabling exploration, ensuring user satisfaction, and facilitating task completion [4].
@@
Compared to traditional search engines that rely on keywords or short phrases, conversational search leverages natural language dialogue for interactions [5, 6, 7, 8, 9], greatly enhancing the efficiency of information exchange and optimizing the user experience. This approach supports more complex user queries, manages longer and more intricate historical contexts, and offers comprehensive capabilities for information integration and processing. Moreover, conversational search enables diverse methods of information delivery and possesses the ability to engage in active interactions with users [10, 11, 12]. For example, these systems can proactively ask clarifying questions [13, 14, 15], make recommendations [16], and guide users in better expressing their needs through dialogue. Fig. 1 illustrates the differences between traditional search and conversational search. In this real example, a traditional search engine (single-shot search) fails to maintain context between queries. After the user queries, “What is information retrieval”, and then follows up with “Tell me some of its famous scholars”, the search engine provides irrelevant results because it does not recognize that the user is still referring to “information retrieval scholars”. In contrast, the conversational search engine (multi-turn search) successfully manages multiple query turns by retaining context, correctly interpreting the follow-up question, and providing accurate information about notable scholars in the field of information retrieval.
@@
Figure 1: Comparison between a traditional search engine and a new conversational search engine on an original and its follow-up query. The term “its” in the second query refers to the term “information retrieval” in the first query. However, it is challenging for traditional search engines to understand such omissions in follow-up user queries.
To conduct a systematic review of the technical advancements enabling this interactive paradigm, we examine four key functional capabilities typically identified in the literature as crucial for conversational search systems, including query reformulation, search clarification, conversational retrieval, and response generation, as illustrated in Fig. 2. These components work collaboratively to manage the multi-turn interactions, understand evolving intent, access information, and present it effectively, ultimately enhancing the overall search experience. It is important to note that recent advancements of powerful LLMs are progressively obscuring the boundaries between these traditionally distinct functions as the LLMs can often perform aspects of reformulation, retrieval, and generation in a more integrated, end-to-end fashion. This survey will explore both the developments within these established functional domains, given their continued relevance in contemporary systems and research, and analyze the implications of emerging integrative trends driven by LLMs with the attendant challenges and opportunities.
@@
Search Clarification. Search clarification is another crucial component in conversational search systems. It enables users to refine their search queries through interactive dialogue, covering scenarios like seeking information, performing tasks, or navigating websites [21, 22, 23]. Users often express their search intents ambiguously or in multifaceted ways. To improve understanding, the system can pose clarifying questions, such as “Do you mean [specific term]?” or “Can you provide more details about [topic]?”, rather than directly answering the query. When the system detects that clarification is needed, it proactively asks these questions to understand the user’s intent better. This approach ensures that the system provides more accurate and relevant search results, enhancing the overall user experience by making the search process more personalized and effective. The primary challenges in search clarification include accurately and efficiently identifying when clarification is necessary and generating appropriate clarifying questions.
@@
Evaluation (Sec 2.3) [ Dataset Analysis: CANARD [73], QReCC [74], TREC-CAsT [75, 76, 77, 78], leaf, text width=27.8em ] [ Evaluation: Lexical Overlapping [73, 74], End-to-End [60, 8], leaf, text width=27.8em ] ] ] [ Search Clarification
@@
Tree-of-clarification [111], CAmbigNQ [22], leaf, text width=27.8em ] ] [ Domain-specific Search
Clarification (Sec 3.4) [ Conversational Recommender System: Conversation Policy [112],
DeepCRS [113], DialogueCRS [114], Knowledge CRS [115],
Topic-guided [116], KG-enhanced [117], SCPR [118], EAR [119],
UNICORN [120], MCMIPL [121], RecLLM [122], iEvalLM [123], leaf, text width=27.8em ] [ Legal: CvsT [51], QGBM [124], LeClari [125], leaf, text width=27.8em ] [ Other: interactive classification [86], image guessing [126],
product description  [127, 128], collaborative building [129],
programming [130, 131], user simulation [132],
social or moral situation [13], task-oriented dialogues [23], math [133],
spatial reasoning [134], collaborative dialogue [135], leaf, text width=27.8em ] ] [ Search Clarification
@@
(Sec 4.2)[ Implicit: ConvDR [144], COTED [26], ZeCo [145], QRACDR [146],
DiSCO [147], leaf, text width=27.8em ] [ Explicit: ConvRelExpand [60], HAConvDR [29], leaf, text width=27.8em ] ] [ Data Augmentation
@@
Results (Sec 5.1)[ Fine-Grained RAG [159], Conv-CoA [143], EvoRAG [160], leaf, text width=27.8em ] ] [ Context Dependency
@@
Attribution (Sec 5.3)[ WebGPT [164], LaMDA [165], WebBrain [166], UniConv [167], Coral [168] , leaf, text width=27.8em ] ] ] [ Domain-specific and
@@
LAPS [179], PersonalizedCIR [180], ConvWebAgent [181], Conv2Q [182],
AdaptivePCIR [183] , leaf, text width=27.8em ] ] ] [ Benchmark and
@@
SIPCIS [195], PACS [196],leaf, text width=27.8em ] [ Generation-based: eRAG [197], PRGR [198], EvalLLMQA [199] ,leaf, text width=27.8em ] [ User-based: URS [200], CCPCS [201], DMCBSDS [202] ,leaf, text width=27.8em ] ] [ Benchmark (Sec 7)[ MSDialog [203], TREC CAsT-19 [75], Qulac [21], CANARD [73],
@@
FOCUSQA [216], TREC iKAT-23 [12], ConvRAG [159], Coral [168]
@@
Our paper differentiates itself from prior surveys on conversational search [4, 220, 221, 222] in two significant ways. First, earlier surveys were conducted prior to the emergence of ChatGPT [^2] ChatGPT: https://openai.com/chatgpt/ and primarily focused on traditional methods without LLMs. LLMs have revolutionized the techniques and application scenarios in this area, leading to the development of real commercial conversational search engines. Our survey incorporates the latest LLM techniques in conversational search and discusses the associated new challenges and opportunities. Second, our survey adopts a more systematic approach. We structure our review from the perspective of information flow in conversational search engines, focusing specifically on the four key modules (i.e., query reformulation, clarification, conversational retrieval, and conversational response generation), while consistently discussing the integrative role and impact of LLMs throughout. A broad overview of conversational search is presented in Fig. 3, which categorizes existing studies using a structured taxonomy.
@@ 2.1 Query Reformulation in Conversational Search
Figure 4: An example illustrating the semantic phenomena of Anaphora and Ellipsis in conversational search. Query reformulation is crucial in addressing these semantic gaps.
In conversational search, query reformulation is critical due to the complex nature of user intent, which is often obscured within follow-up questions and results in context-dependent queries [75, 76, 78]. Unlike traditional information retrieval, conversational search must address two key semantic phenomena: Anaphora and Ellipsis [63, 68]. Anaphora involves expressions of which semantics rely on previous context, while Ellipsis refers to the omission of semantic elements understood from prior context. These phenomena necessitate that current queries are semantically linked to previous historical context, e.g., query-answer pairs, making context tracking essential for accurate responses. Fig. 4 shows a case of query reformulation in conversational search. Adaptive techniques [73] are required to resolve references and fill in omitted information based on conversation history. Consequently, effective query reformulation must be dynamic and context-aware, continuously adapting to the evolving discourse. From the literature, the existing methods can be divided into three categories: 1) Query expansion, 2) Query rewriting, and 3) Hybrid, which are introduced as follows.
@@ 2.1.1 Query Expansion
Relevance term classification transforms query reformulation as a binary classification task where terms in the query context are labeled as relevant or irrelevant. The relevant terms are then appended to the original query as expansion. This approach systematically incorporates important terms, thereby improving the retrieval performance. Typically, Kumar et al. [56] develop a binary term classifier with weak-supervision signals to select relevant terms from the dialog history and pseudo relevance feedback (PRF) to de-contextualize the conversational query. Voskarides et al. [57] propose QuReTeC with distant supervision, whose pseudo labels are readily available and can be inferred from user-system interactions. Lin et al. [58, 228] estimate term importance during query reformulation, which evaluates the significance of different terms within the context, ensuring that the most critical terms are retained or emphasized in the reformulated query. Besides, heuristic methods focus on leveraging linguistic features to reformulate queries, which leverage techniques like part-of-speech tagging, dependency parsing, and co-reference resolution to reconstruct the original query by identifying grammatical structures and resolving references to the previous context. By utilizing syntactic and semantic analysis, these methods [59, 57] enhance the clarity and relevance with the context of the reformulated query and thus improve the search results.
@@ 2.2 Analysis of Existing Datasets
Typical query reformulation datasets designed for conversational search include CANARD [73], QReCC [74], the TREC Conversational Assistance Track (CAsT) series [75, 76, 77, 78] etc. Their construction involves both manual annotation and automatic generation. As aforementioned discussion, the quality of the annotation would greatly affect the search performance. Though with high cost, the human annotators manually rewrite the original queries to provide clear and explicit context, ensuring high-quality annotation. This is done in the TREC Conversational Assistance Track (CAsT) and the recently improved TREC Interactive Knowledge Assistance Track (iKAT) [12, 231] series across the test set, and some of them are exemplified by the CANARD dataset in training data. The rewritten queries provided in the QReCC dataset are inherited from CANARD and expanded automatically by replacing the pronouns with anaphora resolution rules. As a result, its linguistic phenomenon is not as abundant as the fully manually constructed datasets. Besides, the relevance judgments for the documents associated with the original question are determined by whether containing the answers from the annotators, which might not guarantee the corresponding rewritten queries are optimal search queries. Thus, the discrepancy between document assessment and rewritten query annotation for each query turn should be carefully addressed when designing a query reformulation model for downstream search tasks.
@@ 2.4 Limitation and Discussion
Future Paradigm of Query Reformulation. The goal of query reformulation in existing retrieval systems, either ad-hoc or conversational search, is to supply missing search aspects or address ambiguity in the original query. With the development of large language models, some new paradigms of query reformulation have emerged, such as collaboration with different search agents [232], proactively providing reformulated query [233], and reformulating queries with reasoning chains and generative language models [234, 235, 236]. These studies improve query reformulation with various aspects and target different goals without a well-defined paradigm. Furthermore, LLMs might implicitly reformulate queries during the generation process in any conversational systems, or advanced prompting techniques might directly elicit context-aware search behavior. Future work needs to explore the optimal interplay between explicit reformulation modules and the inherent contextual capabilities of LLMs, moving beyond optimizing reformulation in isolation. Thus, it is still unclear what the destination of query reformulation is under the evolved conversational search paradigm, which requires further exploration from the community.
@@ 3 Search Clarification
Conversational search provides users with an interactive interface enabling them to seek information, conduct behavior, or find a website with multi-turn interactions. During the conversation, users may deliver ambiguous, faceted, or unspecific search intents, which should be clarified by the systems for exact query understanding. To this end, the systems are expected to actively ask the users clarifying questions to help them re-articulate their intents. In this section, we summarize studies about asking clarifying questions in conversational search systems. Since the scenarios of different conversational search systems might be various, we first categorize the clarification scenarios as four main parts shown in Fig. 5, including (1) Clarification for conversational document retrieval, (2) Web search clarification, (3) Search clarification for question answering, and (4) Domain-specific search clarification. As for the domain-specific search clarification, we further divide it into different specific domains, including conversational recommender systems, legal search clarification, search clarification for programming, etc. We further summarize studies focusing on search clarification with large language models (LLMs) and finally discuss future directions.
This taxonomy categorization for search clarification scenario elucidates current technical boundaries and identifies promising pathways for developing context-aware clarification architectures, where each of them has distinct objectives, necessitating specialized datasets and learning paradigms. In conversational document retrieval clarification, which requires identifying relevant documents from a corpus, user queries are typically formulated as natural language questions. The primary challenge lies in generating clarifying questions that effectively disambiguate document clusters through strategic partitioning. In contrast, web search clarification focuses on addressing the inherent ambiguity of short queries commonly encountered in commercial search engines, where the challenges involve resolving semantic under-specification within limited lexical context. In the context of question answering systems, which operate in knowledge-intensive environments, the search clarification is crucial for precise factual resolution. Besides, domain-specific search clarification introduces unique complexities from specialized terminologies and vertical knowledge requirements. The inherent differences in operational goals and contextual limitations pose substantial challenges to developing a cohesive framework for cross-scenario clarification systems.
@@ 3.1 Clarification for Conversational Retrieval
Conversational retrieval is a crucial part of a conversational search system. In the retrieval process, search clarification is defined as the multi-turn interaction between the user and the system [21] as shown in Fig. 5 (a). Formally, in each interaction turn $k$, the user inputs a query ($k=1$) or responds to the clarifying question returned by the system ($k>1$). For the system, it asks a clarifying question and retrieves relevant documents using obtained contexts simultaneously in each turn. The fundamental idea of a conversational retrieval system with clarification is to guide users through multiple rounds of queries, prompting them to provide increasingly useful information so the system can return more relevant documents in line with the user’s intentions [237]. In this section, we first summarize the available datasets for clarification in conversational retrieval and introduce clarifying question selection and generation methods. Then we further explore analytical studies that focus on data features instead of proposing new methods. Finally, we introduce some other relevant studies to improve the performance of the systems.